OpenStudy (anonymous):

Find the limit. lim 3 − 3 tan x/sin x − cos x x→π/4

OpenStudy (anonymous):

Could you write out how this looks. the division sign doesn't really indicate where the equation is divided.

OpenStudy (anonymous):

At least put parenthesis where they belong.

OpenStudy (anonymous):

lim (3 − 3 tan x)/(sin x − cos x) x→π/4

OpenStudy (anonymous):

First try to just put in the value pi/4 and decide if the function is defined in that point

OpenStudy (anonymous):

you'll get -3sqrt(2)

OpenStudy (anonymous):

You see that sin pi/4- cos pi/4 equals zero so the function is undefined at pi/4 so what you want to do is to simplify the function

OpenStudy (anonymous):

okaaaaaaaaay thambi

OpenStudy (anonymous):

how do i simplify this?

OpenStudy (chrisplusian):

you can turn your denominator into\[(3-3(\frac{ cosx }{ sinx })\] then split the numerator into two different fractions\[\frac{ 3 }{ sinx-cosx } - \frac{ \frac{ sinx }{ cosx } }{ sinx-cosx }\]

OpenStudy (chrisplusian):

sorry meant numerator

OpenStudy (chrisplusian):

and that should have been three times sine over cosine sorry I am messing it up pretty bad

OpenStudy (anonymous):

loll itss okayy

OpenStudy (anonymous):

i dont get the question

OpenStudy (anonymous):

\[\frac{ 3-3\tan(x) }{ \sin(x) - \cos(x) } \times \frac{ \sin(x) +\cos(x) }{ \sin(x) +\cos(x) }\]

OpenStudy (anonymous):

\[\frac{ 3\sin(x) - 3\frac{ \sin^2(x) }{ \cos(x) } + 3\cos(x) -3\sin(x) }{ \sin^2(x) -\cos^2(x) }\]

OpenStudy (anonymous):

\[\frac{ - 3\frac{ \sin^2(x) }{ \cos(x) } + 3\cos(x) }{ \sin^2(x) -\cos^2(x) }\]

OpenStudy (anonymous):

\[\frac{ - 3\sin^2(x) + 3\cos^2(x) }{ \cos(x)[\sin^2(x) -\cos^2(x)] } = \frac{ -3 }{ \cos(x) }\]

OpenStudy (anonymous):

sorry i was busy

OpenStudy (anonymous):

so i sub in pi/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-3/(1/sqroot2)

OpenStudy (anonymous):

which simplifies into -3sqrt(2)

OpenStudy (anonymous):

-sqroot2/3

OpenStudy (anonymous):

no .. -3 * sqrt(2)

OpenStudy (anonymous):

yeye

OpenStudy (anonymous):

i gett it

OpenStudy (anonymous):

\[-3\sqrt2\]

OpenStudy (anonymous):

like this